# Optimal Best-Arm Identification under Fixed Confidence with Multiple Optima

### v1

@@ VII Sample Complexity Analysis
VII-A Almost-sure Upper Bound
Theorem 6.
Let $\mu$ denote a bandit model from an exponential family. Consider applying Chernoff’s stopping rule with confidence parameter
$\beta(t,\delta)=\log\left(\frac{Ct^{\alpha}}{\delta}\right),$
$\limsup_{\delta\to 0}\frac{\tau_{\delta}}{\log(1/\delta)}\leq T^{*}(\mu),$
where $\tau_{\delta}$ denotes the stopping time.
Proof.
For sufficiently large $t$, i.e., $t\geq t_{0}$, it holds that
$(\hat{\mu}_{1}(t)\geq\hat{\mu}_{a}(t))\vee(\hat{\mu}_{2}(t)\geq\hat{\mu}_{a}(t))\wedge\cdots\wedge(\hat{\mu}_{M}(t)\geq\hat{\mu}_{a}(t))$
for all $a\notin\{1,2,\ldots,M\}$.
Now observe that
$Z(t)$ $\geq\min_{a\notin\{1,\ldots,M\}}Z_{a;1,\ldots,M}(t)$ (78)
$=\min_{a\notin\{1,\ldots,M\}}\sum_{i=1}^{M}N_{i}(t)\,d\left(\hat{\mu}_{i}(t),\hat{\mu}_{a;1,\ldots,M}(t)\right)+N_{a}(t)\,d\left(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,\ldots,M}(t)\right)$ (79)
$=t\cdot\min_{a\notin\{1,\ldots,M\}}\left(\sum_{i=1}^{M}\frac{N_{i}(t)}{t}+\frac{N_{a}(t)}{t}\right)$
$\quad\cdot I_{\frac{N_{1}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t},\ldots,\frac{N_{M}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t}}\left(\hat{\mu}_{1}(t),\ldots,\hat{\mu}_{M}(t),\hat{\mu}_{a}(t)\right).$ (80)
For $a\geq M+1$, the function
$(w,\lambda)\mapsto\left(\sum_{i=1}^{M}w_{i}+w_{a}\right)I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\ldots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\lambda_{1},\ldots,\lambda_{M},\lambda_{a})$
is continuous at $(w^{*}(\mu),\mu)$. Therefore, for any $\varepsilon>0$, there exists $t_{1}\geq t_{0}$ such that for all $t\geq t_{1}$ and all $a\in\{M+1,\ldots,K\}$,
$\left(\sum_{i=1}^{M}\frac{N_{i}(t)}{t}+\frac{N_{a}(t)}{t}\right)$
$\quad\cdot I_{\frac{N_{1}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t},\ldots,\frac{N_{M}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t}}\left(\hat{\mu}_{1}(t),\ldots,\hat{\mu}_{M}(t),\hat{\mu}_{a}(t)\right)$
$\qquad\geq\frac{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}{1+\varepsilon}\cdot I_{\frac{w_{1}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}},\ldots,\frac{w_{M}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}}(\mu_{1},\ldots,\mu_{M},\mu_{a}).$ (81)
Hence, for $t\geq t_{1}$,
$Z(t)$ $\geq\frac{t}{1+\varepsilon}\cdot\min_{a\notin\{1,\ldots,M\}}\left(\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}\right)\cdot I_{\frac{w_{1}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}},\ldots,\frac{w_{M}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}}(\mu_{1},\ldots,\mu_{M},\mu_{a})$
$=\frac{t}{(1+\varepsilon)T^{*}(\mu)}.$ (82)
Therefore,
$\tau_{\delta}$ $=\inf\{t\in\mathbb{N}:Z(t)\geq\beta(t,\delta)\}$ (83)
$\leq t_{1}\vee\inf\left\{t\in\mathbb{N}:\frac{t}{1+\varepsilon}\cdot\frac{1}{T^{*}(\mu)}\geq\log\left(\frac{Ct^{\alpha}}{\delta}\right)\right\},$ (84)
where
$C=K^{M}\left(\frac{4}{M+1}\right)^{(M+1)/2},\quad\alpha=\frac{M+1}{2}.$
@@ VII-A Almost-sure Upper Bound
In the case where $M$ is fixed (not a function of $\delta$), applying the lemma with $c_{1}=\frac{1}{(1+\varepsilon)T^{*}(\mu)}$ and $c_{2}=\frac{C}{\delta}$ yields
@@
$\limsup_{\delta\to 0}\frac{\tau_{\delta}}{\log(1/\delta)}\leq(1+\varepsilon)T^{*}(\mu).$
@@ VII-B Asymptotic Optimality in Expectation
Let $\mu$ be a bandit model from an exponential family. Using Chernoff’s stopping rule with
@@
$\limsup_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau_{\delta}]}{\log(1/\delta)}\leq T^{*}(\mu).$ (85)
@@
Following a similar strategy as in (Garivier2016OptimalBA, , Theorem 14), we simplify notation by assuming that the bandit means satisfy
$\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{M}>\mu_{M+1}\geq\mu_{M+2}\geq\cdots\geq\mu_{K}.$
@@
On the event $\mathcal{E}_{T}$, for $t\geq h(T)$ and any $a\notin\{1,2,\dots,M\}$, we have $\hat{\mu}_{i}(t)\geq\hat{\mu}_{a}(t)$ for all $i\in\{1,\dots,M\}$, and the Chernoff-type statistic satisfies
@@
$\qquad\geq\min_{a\notin\{1,\dots,M\}}Z_{a;1,\dots,M}(t)$ (86)
$\qquad=\min_{a\notin\{1,\dots,M\}}\left[N_{a}(t)d(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,\dots,M}(t))+\sum_{i=1}^{M}N_{i}(t)d(\hat{\mu}_{i}(t),\hat{\mu}_{a;1,\dots,M}(t))\right]$ (87)
$\qquad=t\cdot g\left(\hat{\mu}(t),\left\{\frac{N_{i}(t)}{t}\right\}_{i=1}^{K}\right),$ (88)
@@
$g(\mu^{\prime},w^{\prime}):=\min_{a\notin\{1,\dots,M\}}\left(\sum_{i=1}^{M}w^{\prime}_{i}+w^{\prime}_{a}\right)I_{\left\{\frac{w^{\prime}_{i}}{\sum_{j=1}^{M}w^{\prime}_{j}+w^{\prime}_{a}}\right\}_{i=1}^{M}}(\mu^{\prime}_{1},\dots,\mu^{\prime}_{M},\mu^{\prime}_{a}).$
Using (Garivier2016OptimalBA, , Lemma 20), for $T\geq T_{\varepsilon}$, define
@@
Then, for $T\geq T_{\varepsilon}$, on $\mathcal{E}_{T}$,
$\min(\tau_{\delta},T)$ $\leq\sqrt{T}+\sum_{t=\sqrt{T}}^{T}\mathbbm{1}_{\tau_{\delta}>t}$ (89)
$\leq\sqrt{T}+\sum_{t=\sqrt{T}}^{T}\mathbbm{1}\left\{\max_{b_{1},\dots,b_{M}}\min_{a\notin\{b_{1},\dots,b_{M}\}}Z_{a;b_{1},\dots,b_{M}}(t)\leq\beta(t,\delta)\right\}$ (90)
@@
Then, for any $T\geq\max(T_{0}(\delta),T_{\varepsilon})$,
@@
$\mathbb{E}_{\mu}[\tau_{\delta}]\leq T_{0}(\delta)+T_{\varepsilon}+\sum_{T=1}^{\infty}BT\exp(-CT^{1/8}).$
@@
$T_{0}(\delta)$ $\leq C(\eta)+\inf\left\{T\in\mathbb{N}:\frac{1}{C^{*}_{\varepsilon}(\mu)}\log\left(\frac{CT^{\alpha}}{\delta}\right)\leq\frac{T}{1+\eta}\right\}$ (93)
$\leq C(\eta)+\inf\left\{T\in\mathbb{N}:\frac{C^{*}_{\varepsilon}(\mu)}{1+\eta}T\geq\log\left(\frac{CT^{\alpha}}{\delta}\right)\right\}.$ (94)
@@
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau_{\delta}]}{\log\frac{1}{\delta}}\leq\frac{1+\eta}{C^{*}_{\varepsilon}(\mu)}.$
@@
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau_{\delta}]}{\log\frac{1}{\delta}}\leq T^{*}(\mu).$

### v2

@@ VII Sample Complexity Analysis
In this section, we analyze the sample complexity of the modified C-Tracking and D-Tracking algorithms introduced in Section V. We begin by establishing the following result.
@@
Based on Lemma 7, the following subsections derive upper bounds on the sample complexity, which asymptotically match the lower bound established in Theorem 5.
VII-A Almost-sure Upper Bound
Theorem 8.
Let $\mu$ denote a bandit model from an exponential family. Define
$C=K^{M}\left(\frac{4}{M+1}\right)^{(M+1)/2},\quad\alpha=\frac{M+1}{2},$
Consider applying Chernoff’s stopping rule with confidence parameter
$\beta(t,\delta)=\log\left(\frac{Ct^{\alpha}}{\delta}\right),$
$\limsup_{\delta\to 0}\frac{\tau}{\log(1/\delta)}\leq T^{*}(\mu),$
where $\tau$ denotes the stopping time, and $T^{*}(\mu)$ is defined in Theorem 5.
Proof.
For sufficiently large $t$, i.e., $t\geq t_{0}$, it holds that
$(\hat{\mu}_{1}(t)\geq\hat{\mu}_{a}(t))\vee(\hat{\mu}_{2}(t)\geq\hat{\mu}_{a}(t))\wedge\cdots\wedge(\hat{\mu}_{M}(t)\geq\hat{\mu}_{a}(t))$ (105)
for all $a\notin[M]$.
In addition, by using the C-Tracking and D-Tracking sampling rule, for $t_{1}$ sufficiently large ($t_{1}>t_{0}$), for any $t\geq t_{1}$ and any $a\notin[M]$, we also have
$\max_{a\notin[M]}S_{a;1,2,\cdots,M}(t)$
$\quad=\max_{a\notin[M]}\frac{N_{a}(t)}{t}\big[d(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,2,\cdots,M}(t)+\zeta(t))$
$\qquad-d(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,2,\cdots,M}(t))\big]$
$\qquad-\min_{i\in[M]}\frac{N_{i}(t)}{t}d(\hat{\mu}_{i}(t),\hat{\mu}_{a;1,2,\cdots,M}(t))$
$\quad\leq-\frac{1}{2}\min_{a\notin[M]}\min_{i\in[M]}w_{i}^{*}d\bigg(\mu_{i},\frac{\sum_{i=1}^{M}w_{i}^{*}\mu_{i}+w_{a}^{*}\mu_{a}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}\bigg)<0.$ (106)
Now for $t>t_{1}$, observe that
$Z(t)$ $\geq\min_{a\notin[M]}Z_{a;1,\ldots,M}(t)$
$=\min_{a\notin[M]}\sum_{i=1}^{M}N_{i}(t)\,d\left(\hat{\mu}_{i}(t),\hat{\mu}_{a;1,\ldots,M}(t)\right)$
$\quad+N_{a}(t)\,d\left(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,\ldots,M}(t)\right)$
$=t\min_{a\notin[M]}\left(\sum_{i=1}^{M}\frac{N_{i}(t)}{t}+\frac{N_{a}(t)}{t}\right)$
$\quad\times I_{\frac{N_{1}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t},\ldots,\frac{N_{M}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t}}$
$\qquad\left(\hat{\mu}_{1}(t),\ldots,\hat{\mu}_{M}(t),\hat{\mu}_{a}(t)\right).$ (107)
For $a\geq M+1$, the function
$(w,\lambda)\mapsto\left(\sum_{i=1}^{M}w_{i}+w_{a}\right)$
$\qquad\times I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\ldots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\lambda_{1},\ldots,\lambda_{M},\lambda_{a})$
is continuous at $(w^{*}(\mu),\mu)$. Therefore, for any $\varepsilon>0$, there exists $t_{2}\geq t_{1}$ such that for all $t\geq t_{2}$ and all $a\in\{M+1,\ldots,K\}$,
$\left(\sum_{i=1}^{M}\frac{N_{i}(t)}{t}+\frac{N_{a}(t)}{t}\right)$
$\qquad\times I_{\frac{N_{1}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t},\ldots,\frac{N_{M}(t)/t}{\sum_{i=1}^{M}N_{i}(t)/t+N_{a}(t)/t}}$
$\qquad\left(\hat{\mu}_{1}(t),\ldots,\hat{\mu}_{M}(t),\hat{\mu}_{a}(t)\right)$
$\quad\geq\frac{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}{1+\varepsilon}$
$\qquad\times I_{\frac{w_{1}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}},\ldots,\frac{w_{M}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}}(\mu_{1},\ldots,\mu_{M},\mu_{a}).$ (108)
Hence, for $t\geq t_{2}$,
$Z(t)$ $\geq\frac{t}{1+\varepsilon}\cdot\min_{a\notin[M]}\left(\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}\right)$
$\qquad\times I_{\frac{w_{1}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}},\ldots,\frac{w_{M}^{*}}{\sum_{i=1}^{M}w_{i}^{*}+w_{a}^{*}}}(\mu_{1},\ldots,\mu_{M},\mu_{a})$
$=\frac{t}{(1+\varepsilon)T^{*}(\mu)}.$ (109)
Therefore,
$\tau$ $=\inf\{t\in\mathbb{N}:Z(t)\geq\beta(t,\delta)\}$
$\leq t_{2}\vee\inf\left\{t\in\mathbb{N}:\frac{t}{1+\varepsilon}\cdot\frac{1}{T^{*}(\mu)}\geq\log\left(\frac{Ct^{\alpha}}{\delta}\right)\right\},$ (110)
where
$C=K^{M}\left(\frac{4}{M+1}\right)^{(M+1)/2},\quad\alpha=\frac{M+1}{2}.$
In the case where $M$ is fixed (not a function of $\delta$), applying the Lemma 7 with $c_{1}=\frac{1}{(1+\varepsilon)T^{*}(\mu)}$ and $c_{2}=\frac{C}{\delta}$ yields
@@ VII-A Almost-sure Upper Bound
$\limsup_{\delta\to 0}\frac{\tau}{\log(1/\delta)}\leq(1+\varepsilon)T^{*}(\mu).$
@@ VII-B Asymptotic Optimality in Expectation
Let $\mu$ be a bandit model from an exponential family. Define
$C=K^{M}\left(\frac{4}{M+1}\right)^{(M+1)/2},\quad\alpha=\frac{M+1}{2},$
Using Chernoff’s stopping rule with
@@
$\limsup_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau_{\delta}]}{\log(1/\delta)}\leq T^{*}(\mu),$ (111)
where $T^{*}(\mu)$ is defined in Theorem 5.
Remark 10.
Some remarks are in order.
  • Consider the special Gaussian bandit $\mathcal{N}(\mu,\sigma^{2})$ where $\mu_{1}=\mu_{2}=\cdots=\mu_{M}=\Delta$ and $\mu_{M+1}=\mu_{M+2}=\cdots=\mu_{K}=0$ for some $\Delta>0$. Then, for any $a\notin[M]$ we have
    $I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\cdots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\mu_{1},\cdots,\mu_{M},\mu_{a})$
    $\quad=\bigg[\frac{\beta_{a}(w)(1-\beta_{a}(w))^{2}}{2\sigma^{2}}\Delta^{2}+\frac{(1-\beta_{a}(w))\beta_{a}^{2}(w)}{2\sigma^{2}}\Delta^{2}\bigg]$
    $\quad=\frac{\beta_{a}(w)(1-\beta_{a}(w))}{2\sigma^{2}}\Delta^{2},$ (112)
    where
    $\beta_{a}=\frac{\sum_{i=1}^{M}w_{i}}{\sum_{i=1}^{M}w_{i}+w_{a}}.$ (113)
    Hence, by (19) it holds that
    $T^{*}(\mu)$ $=\frac{2\sigma^{2}}{\sup_{w\in\Sigma_{K}}\min_{a\notin[M]}\beta_{a}(w)(1-\beta_{a}(w))\Delta^{2}}$
    $=\Theta\bigg(\frac{1}{\Delta^{2}}\bigg).$ (114)
    This implies that the expected sample complexity diverges to infinity as $\Delta\to 0$ for $\Delta>0$. When $\Delta=0$ and $M=K$ is known, any arm can be selected as mentioned above, and hence the sample complexity is finite (in fact, zero).
  • For each bandit instance $\lambda\in\mathcal{S}$, let $i^{*}(\lambda)$ denote the set of all optimal arms. Further, for each $i\in\{1,2,\dots,M\}$, define
    $\neg i:=\{\lambda\in\mathcal{S}:i\notin i^{*}(\lambda)\}=\bigcup_{a\neq i}\{\lambda\in\mathcal{S}:\lambda_{a}>\lambda_{i}\}.$
    It is straightforward to verify that $\neg i\supset\text{Alt}(\mu)$ for all $i$. Therefore, by [1, Theorem 1] and (18), we have
    $T^{*}(\mu)\leq D(\mu)^{-1},$
    where
    $D(\mu)=\max_{i\in i^{*}(\mu)}\max_{w\in\Sigma_{K}}\inf_{\lambda\in\neg i}\sum_{a=1}^{K}w_{a}\,d(\mu_{a},\lambda_{a}).$ (115)
    Here, $D(\mu)^{-1}$ corresponds to the sample complexity of the same bandit problem in the unknown-$M$ setting [1]. This highlights that knowing $M$ can improve the sample complexity. Technically, the stopping time based on the generalized log-likelihood ratio (GLLR) statistic $Z_{a;b_{1},b_{2},\dots,b_{M}}(t)$ in (45) is only well-defined when $M$ is known (see Section V).
@@
Following a similar strategy as in [3, Theorem 14], we simplify notation by assuming that the bandit means satisfy
$\mu_{1}=\mu_{2}=\cdots=\mu_{M}>\mu_{M+1}\geq\mu_{M+2}\geq\cdots\geq\mu_{K}.$
@@
In addition, by Lemma [3, Lemma 20], for any $\varepsilon>0$ there exists a constant $T_{\varepsilon}$ such that for $T\geq T_{\varepsilon}$, it holds on $\mathcal{E}_{T}$, for either C-Tracking or D-Tracking,
$\max_{a}\bigg|\frac{N_{a}(t)}{t}-w_{a}^{*}(\mu)\bigg|\leq 3(K-1)\varepsilon,\qquad\forall t\geq\sqrt{T}.$ (116)
Hence, by using the C-Tracking and D-Tracking sampling rule, for any sufficient small $\varepsilon$, there exists $T^{*}_{\varepsilon}>T_{\varepsilon}$ such that for any $T>T_{\varepsilon}^{*}$ we have
$\sup_{a\notin[M]}S_{a;1,2,\cdots,M}(t)$
$\quad=\max_{a\notin[M]}\frac{N_{a}(t)}{t}\big[d(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,2,\cdots,M}(t)+\zeta(t))$
$\qquad-d(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,2,\cdots,M}(t))\big]$
$\quad\qquad-\min_{i\in[M]}\frac{N_{i}(t)}{t}d(\hat{\mu}_{i}(t),\hat{\mu}_{a;1,2,\cdots,M}(t))$
$\quad\leq 0,\qquad\forall t\geq\sqrt{T}.$ (117)
On the event $\mathcal{E}_{T}$, for $t\geq h(T)$ and any $a\notin[M]$, we have $\hat{\mu}_{i}(t)\geq\hat{\mu}_{a}(t)$ for all $i\in[M]$ and $S(t)\leq 0$, and the Chernoff-type statistic satisfies
@@
$\quad\geq\min_{a\notin[M]}Z_{a;1,\dots,M}(t)$ (118)
$\quad=\min_{a\notin[M]}\big[N_{a}(t)d(\hat{\mu}_{a}(t),\hat{\mu}_{a;1,\dots,M}(t))$
$\qquad+\sum_{i=1}^{M}N_{i}(t)d(\hat{\mu}_{i}(t),\hat{\mu}_{a;1,\dots,M}(t))\big]$ (119)
$\quad=t\cdot g\left(\hat{\mu}(t),\left\{\frac{N_{i}(t)}{t}\right\}_{i=1}^{K}\right),$ (120)
@@
$g(\mu^{\prime},w^{\prime})$ $:=\min_{a\notin[M]}\left(\sum_{i=1}^{M}w^{\prime}_{i}+w^{\prime}_{a}\right)$
$\qquad\times I_{\left\{\frac{w^{\prime}_{i}}{\sum_{j=1}^{M}w^{\prime}_{j}+w^{\prime}_{a}}\right\}_{i=1}^{M}}(\mu^{\prime}_{1},\dots,\mu^{\prime}_{M},\mu^{\prime}_{a}).$ (121)
Using [3, Lemma 20], for $T\geq T_{\varepsilon}^{*}$, define
@@
Then, for $T\geq T_{\varepsilon}^{*}$, on $\mathcal{E}_{T}$,
$\min(\tau,T)$
$\leq\sqrt{T}+\sum_{t=\sqrt{T}}^{T}\mathbbm{1}_{\tau>t}$ (122)
$\leq\sqrt{T}$
$\quad+\sum_{t=\sqrt{T}}^{T}\mathbbm{1}\left\{\max_{b_{1},\dots,b_{M}}\min_{a\notin\{b_{1},\dots,b_{M}\}}Z_{a;b_{1},\dots,b_{M}}(t)\leq\beta(t,\delta)\right\}$ (123)
@@
Then, for any $T\geq\max(T_{0}(\delta),T_{\varepsilon}^{*})$,
@@
$\mathbb{E}_{\mu}[\tau]\leq T_{0}(\delta)+T_{\varepsilon}^{*}+\sum_{T=1}^{\infty}BT\exp(-CT^{1/8}).$
@@
$T_{0}(\delta)$
$\quad\leq C(\eta)+\inf\left\{T\in\mathbb{N}:\frac{1}{C^{*}_{\varepsilon}(\mu)}\log\left(\frac{CT^{\alpha}}{\delta}\right)\leq\frac{T}{1+\eta}\right\}$ (126)
$\quad\leq C(\eta)+\inf\left\{T\in\mathbb{N}:\frac{C^{*}_{\varepsilon}(\mu)}{1+\eta}T\geq\log\left(\frac{CT^{\alpha}}{\delta}\right)\right\}.$ (127)
@@
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau]}{\log\frac{1}{\delta}}\leq\frac{1+\eta}{C^{*}_{\varepsilon}(\mu)}.$
@@
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau]}{\log\frac{1}{\delta}}\leq T^{*}(\mu).$